OpenStudy (anonymous):

find the slope of the line passing through the given point E(-1,2) F(5,0) and show your work. please help :(

OpenStudy (tkhunny):

Well, please show your work. How does one calculate slope, given two points?

OpenStudy (anonymous):

slope is calculated by the formula: \[\frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\]

OpenStudy (anonymous):

you have 2 points. select one point for (x1,y1, and the other for (x2,y2).

OpenStudy (anonymous):

then plug your numbers into the formula. does this make sense?

OpenStudy (anonymous):

i did and i got 1/2 but my teacher said it was wrong

OpenStudy (anonymous):

i still cnt find the answer

OpenStudy (tkhunny):

5 - (-1) = 5 + 1 = 6 Not 4

OpenStudy (anonymous):

thanks... do u think you could help with another one

OpenStudy (tkhunny):

Only if you post it. A different thread is preferred.